# Large Language Models and Games: A Survey and Roadmap

### v1

@@ 4 A Roadmap for Future Applications of LLMs in Games
The previous section attempted to group current research in LLMs and games into a typology focused on the roles an LLM is asked to play. As part of this exercise, we identified a number of roles that have been heavily researched. Unsurprisingly, the role of player and automated designer have received attention: this matches the general trends within AI and Games research more broadly [5]. Following general trends in Game AI for playing or generating content, LLM-based approaches are likely to flourish via community events, benchmarks and competitions, with first steps already being taken in this direction [77]. Based on the roles listed in Section 3, we identify below some gaps found in the literature, and lay out possible research directions that leverage the power of LLMs in new ways.
While academic interest in design assistance within games has blossomed in the last decade, we find that the potential of LLMs has so far been underutilized. LLM design assistants either ask too much of a human designer in terms of creative interpretation and actual development [74] or too little, demoting them to content curator [75]. Past research in mixed-initiative systems [68] assumes a more co-creative initiative from both human and machine, and the power of LLMs as conversational agents matches the original vision of a creative dialogue between initiatives [70]. Therefore, a promising unexplored direction lies in a more procedural assistance (see Section 3.8) where the LLM not only produces output but also reasons about it to the human designer. LLMs seem especially well-suited for this task, as the context is retained and the designer can iteratively refine past products that the LLM has generated. However, concerns of LLMs’ limited memory may arise (see Section 5) in long-term design process. On the other hand, iterative refining is not as straightforward for other state-of-the-art technology such as LMMs, despite some promising results via e.g. InstructPix2Pix [78]. It is expected that such applications will raise new challenges in terms of hallucinations, explainability [79], capturing or modeling designer intent [71], and more. We discuss such challenges further in Section 6.
@@
Another role seemingly well-suited for LLMs that has received limited attention is that of commentator or reteller. Work so far has focused on automating the commentary of streamers or eSport casters [42]. While this direction is still largely uncharted, there are more directions that could leverage LLMs for streamer assistance rather than automation (and replacement). Rather than narrate events occurring within the game (or video stream), LLMs can summarize the audience interactions and engagement levels—thus acting as a commentator not of the game but of the audience watching it. This could allow a human streamer to better keep track of topics discussed in the chat, and engage as needed without having to read every comment. While this has been identified as a research direction for AI already [80], it has not yet been implemented. Under the role of streamer assistance, issues of explainability of the LLM’s commentary would become pertinent (e.g. to address one audience member by name); we revisit this in Section 5.
@@
Finally, it is worth noting that one pillar of AI in games research is missing altogether from LLM research and applications surveyed: that of modeling players [5]. LLMs appear to offer promising, yet admittedly unexplored, methods for affective game computing [81]. At first glance, this is not particularly surprising: while player modeling often relies on supervised machine learning algorithms, it is not as clear how this can be achieved via text input or output. In principle, an LLM could predict affective state transitions such as “the game is more engaging now” and thereby adapt the game environment to elicit a supposedly more engaging experience for the player. Learning such transitions builds on the experience-driven procedural content generation paradigm [82] but with an LLM acting as the player experience model. We therefore envision the fine tuning of LLMs so that they learn to represent and infer player experience transitions based on in-game observations and demonstrations of experience. As highlighted in Section 5, however, current LLMs struggle to capture user intent during conversation—let alone more ill-defined concepts such as players’ emotion or engagement [83]. Current datasets on affect in games are formatted as continuous or categorical variables, often fluctuating over time [84], which would be challenging to format as text without processing. While perhaps using language as input or output for the player model requires some innovative pre-processing or more advanced LLM technologies, the underlying transformer architecture and attention-based algorithms show promise already. We expect more research on player modeling powered by transformers, such as leveraging behavior transformers [85] to imitate human playtraces grouped by playstyle [86].
@@
To wrap up, we believe that every role an LLM could be called to play in (or around) a game identified in Section 3 could benefit from additional attention. This technology remains nascent, and changes are forthcoming which may address several limitations we identified both in the above paragraphs and, more extensively, in Section 5. The natural language capabilities (especially for text generation) make LLMs ideal conversational assistants (for a player, a designer, a game master, or a streamer). The ability of LLMs to consume and reason from text corpora also opens new possibilities for automated design moving beyond tile-based level generation (which needs carefully crafted corpora) and more towards open-ended content such as game narratives [87, 88, 89, 90] or even game design documents. The potential of LLMs in that regard is already voiced by many evangelists in the field, but research on actual implementation of such ideas and on addressing the IP concerns they may raise (see Section 6) are still forthcoming.
@@ 5 Limitations of LLMs in games
Large language models have exciting potential for video games, but they also come with inherent limitations. Mainly, LLMs suffer from hallucinations [32, 91], meaning that they will output plausible but false statements simply because they are a probable sequence of words. Hallucinations are inevitable given how the world is described to the machine [92]; LLMs lack grounding, so the text they generate is detached from constraints of reality. Yet LLMs always “act” confidently in their responses, even when wholly mistaken. Beyond hallucinations, LLMs suffer from factual errors [93, 94, 95], outputting responses that are wrong even though the LLM has access to information that proves otherwise. In the context of video games, these limitations affect certain applications of LLMs more than others, for example NPCs may hallucinate quests that do not exist in the game, or a player assistant may provide suggestions to the user based on wrong assumptions.
@@
Another limitation of using LLMs in video games is that LLMs sometimes struggle to capture user intent. This is especially evident with expressions of sarcasm [96]. The ability to capture user intent is important for applications of LLMs that converse directly with the player. An LLM that correctly understands user intent then also correctly understands the context of the conversation, which is not always the case for current LLMs. Many LLMs make errors in correctly understanding user requests [97], and clarifying to the LLM multiple times results in a frustrating experience for the user. This limitation is most relevant to cases where the LLM is in direct conversation with the user, e.g. as design assistant, player assistant or game master. Depending on how much the LLM output controls the user experience (e.g. as game master or offering production assistance to a human designer), the inability to capture user intent can be a source of frustration.
@@
On a larger scale, LLMs suffer from losing context and struggling with continuity. This is because the “memory” of an LLM is constrained by its context size, which limits the extent of its inputs and outputs, as well as its response time due to the attention mechanism [10]. The longer the conversation, the less likely it is that the LLM will recall early events [98]. In video games, it is possible to separately summarize the game events (see Section 3.4) and process them as part of the input to the LLM. As a game progresses past a few game sessions, however, this summary may still be too long or details of increasing significance will be omitted, leading to a degraded performance. This is especially relevant for roles requiring a long-term engagement, such as LLM-powered retellers or game masters. In Infinite Craft (see Section 3.6), this is handled by an external database that stores and lookups past combination rules—ensuring consistency in future uses of the same mechanic. However, LLMs could theoretically tackle this issue directly. A Retrieval-Augmented Generation (RAG) system [99] could address this limitation, drawing from a database containing vector representations or other latent representations of pertinent text or data. When the text generator processes a sequence, the RAG system would retrieve similar entries from this external data source. This would hypothetically provide a streamlined archive of game events and actions for the LLMs to consult in order to generate a consistent narrative progression.
@@
Finally, the implementation and deployment of LLMs to video game applications is still very limited. A video game is a domain where responsiveness is vital for players, so it follows that LLMs should also be able to provide their responses quickly. Unfortunately, while research on more efficient and faster architectures is being carried out [100], the real-time application of LLMs is still not plausible. This is especially evident in other domains such as design applications, where “real time” responses are generated in around 30 seconds to over a minute [101].
@@ 6 Ethical Issues with LLMs in games
With the improvement of AI methods applied to video games over the recent years, many questions regarding their ethics and real-world impact have been raised [102]. Using LLMs raises ethical issues regarding sustainability, copyright, explainability, and biases. Naturally, each of these issues has serious implications in the field of video games.
@@
The reliance of LLMs on training data and training time raises concerns regarding their carbon footprint. Beyond training costs, inference over the model’s lifespan has a greater environmental impact due to constant querying [103, 104]. Factors like renewable and local energy, better model architectures, and more meaningful (and thus less wasteful) training data can mitigate this. In the context of LLMs for digital games, sustainability remains crucial, considering the carbon footprint of frequent queries during gameplay (e.g. for Game Master or NPC responses, or for LLM-powered players). This is especially pertinent if the LLM is intended to run locally, on consumer-level hardware which is usually powered by non-renewable sources.
@@
When it comes to copyright, issues apply to the input data, the output data, and the model itself. LLMs trained on data under copyright is an unfortunate common practice [105], deservedly raising public outrage [106, 107]. The models themselves have different copyright licenses applied, which can also lead to artifacts they generate to fall under the public domain [11, 108]. For the game industry, matters of IP and copyright are extremely important. This is as much a concern regarding having the company’s copyrighted content somehow used as training by competitors, as it is about using LLMs that can produce material that the company cannot copyright. It is important to note here that, at least when it comes to the latter concern, the role the LLM takes is very pertinent. If an LLM or LMM produces content automatically (see Section 3.7), past legal consensus in the US indicates that the material can not be copyrighted [109]. If an LLM or LMM acts as an “assistive tool” [110] to a designer (especially for conceptual assistance, see Section 3.8) then the extensive and impactful human effort needed to transform these concepts into game design and game art likely makes the final product copyrightable [110]. The limited rulings in copyright courts regarding this, however, and the “likely” caveat we include in our own text, understandably would make game companies hesitant to tread in untested waters for major game IPs beyond e.g. small-scale indie productions [45]. For researchers, however, the ethical issues of copyright breach and exploitation by large corporations, and the public outcry for the above, leave a bad taste and make research in LLMs less palatable [111].
@@
In applications, understanding how a final result or product is reached is a extremely crucial, particularly when a product is iteratively refined as with design assistants (see Section 3.8). This is a problem of explainability [79], whereas LLMs are inherently opaque in their generation process. In [112], the authors highlight different methods to improve the explainability of language models, such as concept-based explanations or saliency maps. Particularly for LLMs, the self-explanation applied via the chain-of-thought (CoT) [113] reasoning has received attention by the research community [114, 115]. While this method adds a layer of explained reasoning to the generated output, there are multiple examples in the literature that demonstrate how this reasoning may just be an illusion of reasoning capabilities. Such examples include disregarding the provided reasoning in the final output [116], or reaching the correct solution via incorrect steps in math problems [117]. In video-games, explainability is paramount across roles, ensuring gameplay coherence and user engagement.
@@
Finally, biases emerge as LLMs are trained on a large corpus, usually scraped from the (Western-focused part of the) internet. This allows models to capture a current reality snapshot, which is advantageous for a conversational or questions answering model, though it requires curating this data from different kinds of biases. Some biases, such as social stereotypes, could be targeted and alleviated; others, such as exclusionary norms, pose greater challenges. In video games, we identify two main concerns when interacting with an LLM: toxic behavior, and stereotypes or incorrect notions. Toxic behavior is a harmful property that a language model may learn from its training corpus, which often contains text from community-based fora or social platforms. Tools that combat toxic language in video games are constantly evolving, with some even blocking chat messages before they are even delivered to the user [118, 119]. Therefore, similar applications could theoretically be developed to target toxic outputs from language models. Unlike human players, however, when an LLM plays the role of an NPC, it should align with the game themes and avoid any kind of toxic language or racial slurs. This requires developers to ensure proper behavior of the model through data cleaning, if the model is trained from scratch, or supplying tailored data if finetuning it to their needs. Addressing prejudices such as stereotypes and incorrect notions is complex, as they are not necessarily related to single words or expressions, but instead present themselves as a collection of ideals that can be wrong at best, and harmful at worst. An NPC LLM may exhibit real world stereotypes that can impact negatively the player experience, although we argue that the impact of prejudices from a player commentator or game master is much stronger and disturbing due to their perceived authority.

### v2

@@ 4 A Roadmap for Future Applications of LLMs in Games
The previous section attempted to group current research in LLMs and games into a typology focused on the roles an LLM is asked to play. As part of this exercise, we identified a number of roles that have been heavily researched. Unsurprisingly, the role of player and automated designer have received most attention: this matches the general trends within AI and games research more broadly [7]. Following general trends in Game AI for playing or generating content, LLM-based approaches are likely to flourish via community events, benchmarks and competitions, with first steps already being taken in this direction [121]. Based on the roles listed in Section 3, we identify below some gaps found in the literature, and lay out possible research directions that leverage the power of LLMs in new ways.
While academic interest in design assistance within games has blossomed in the last decade, we find that the potential of LLMs has so far been underutilized. LLM design assistants either ask too much of a human designer in terms of creative interpretation and actual development [112] or too little, demoting them to content curator [115]. Past research in mixed-initiative systems [103] assumes a more co-creative initiative from both human and machine, and the power of LLMs as conversational agents matches the original vision of a creative dialogue between initiatives [108]. Therefore, a promising unexplored direction lies in a more procedural assistance (see Section 3.9) where the LLM not only produces output but also reasons about it to the human designer. LLMs seem especially well-suited for this task, as the context is retained and the designer can iteratively refine past products that the LLM has generated. However, concerns of LLMs’ limited memory may arise (see Section 5) in long-term design processes. On the other hand, iterative refining is not as straightforward for other state-of-the-art technologies such as LMMs, despite some promising results via e.g. InstructPix2Pix [122]. It is expected that such applications will raise new challenges in terms of hallucinations, explainability [123], capturing or modeling designer intent [109], and more. We discuss such challenges further in Section 6.
@@
Another role seemingly well-suited for LLMs that has received limited attention is that of commentator or reteller. Work so far has focused on automating the commentary of streamers or eSport casters [59]. While this direction is still largely uncharted, there are more directions that could leverage LLMs for streamer assistance rather than automation (and replacement). Rather than narrate events occurring within the game (or video stream), LLMs can summarize the audience interactions and engagement levels —thus acting as a commentator not of the game but of the audience watching it. This could allow a human streamer to better keep track of topics discussed in the chat, and engage as needed without having to read every comment. While this has been identified as a research direction for AI already [124], it has yet to be implemented. Under the role of streamer assistance, issues of explainability of the LLM’s commentary would become pertinent (e.g. to address one audience member by name); we revisit this in Section 5. It is worth noting that streamers have already begun to explore AI assistance to their streams. To the best of our knowledge, the most relevant example is YouTube user Criken who plays alongside an AI assistant as a conversational agent, having a dialogue with an otherwise out-of-the-box LLM either embedded within the game [^12] https://youtu.be/dQ-7-r5aM1U?feature=shared or as part of the stream [^13] https://youtu.be/KhE9NhUqtBc?feature=shared. This example is not explicitly targeted for reacting to in-game events or audience interaction (and acts more as NPC than commentator) but indicates that some streamers are open to the use of such technology for their craft.
@@
Despite a few attempts to leverage LLMs for games user research (see Section 3.5), there is much unexplored potential in this direction. LLMs so far are used to cluster gameplay logs [66, 70], but they could also explain the groupings in the form of e.g. play personas [125, 68] described in natural language. Such a task would raise issues of explainability and privacy more broadly (see Section 5), and would likely still involve a game designer or user researcher in the loop to ensure the quality of conclusions drawn from this process. More importantly, moving from these log-based clusters towards capturing the player’s experience or emotional state [126] remains an open challenge. In principle, an LLM could predict affective state transitions such as “the game is more engaging now” and thereby adapt the game environment to elicit a supposedly more engaging experience for the player. Learning such transitions builds on the experience-driven procedural content generation paradigm [127] but with an LLM acting as the player experience model. Future research could explore how LLMs can be fine-tuned to represent and infer player experience transitions based on in-game observations and demonstrations of experience. Two challenges need to be addressed for this: (a) representations of game states as natural language and (b) hallucinations of human experience. For the former, potentially leveraging work of LLM players (see Section 3.1) and how they pass the game state to the LLM seems a promising first step. For the latter challenge, however, current LLMs struggle to capture user intent during conversation —let alone more ill-defined concepts such as players’ emotion or engagement [128]. Current datasets on affect in games are formatted as continuous or categorical variables, often fluctuating over time [129], which would be challenging to format as text without processing. While perhaps using language as input or output for the player model requires some innovative pre-processing or more advanced LLM technologies, the underlying GPT architecture shows promise already. Broekens et al. showed that ChatGPT could detect emotion in English text [130], although admittedly games include many more modalities (e.g. visuals, audio) than pure text, which is mainly relegated to narrative [131]. We expect more research on player modeling powered by transformers, if not LLMs directly, such as leveraging behavior transformers [132] to imitate human playtraces grouped by playstyle [133].
@@
To wrap up, we believe that every role an LLM could be called to play in (or around) a game identified in Section 3 could benefit from additional attention. This technology remains nascent, and changes are forthcoming which may address several limitations we identified both in the above paragraphs and, more extensively, in Section 5. The natural language capabilities (especially for text generation) make LLMs ideal conversational assistants (for a player, a designer, a Game Master, or a streamer). The ability of LLMs to consume and reason from text corpora also opens new possibilities for automated design moving beyond tile-based level generation (which needs carefully crafted corpora) and more towards open-ended content such as game narratives [134, 135, 136, 137, 138, 139] or even game design documents. The potential of LLMs in that regard is already voiced by many evangelists in the field, but research on actual implementation of such ideas and on addressing the IP concerns they may raise (see Section 6) are still forthcoming.
While the focus of this paper is on what LLMs can do for games, we do not underestimate what games can do for LLMs. One of the watershed moments for AI and games research was the article by Laird and Van Lent naming games as the “killer app” for human-level AI [140]. This remains true for LLMs today: games are ideally poised for LLM research. Not only do games produce rich multimodal data (ideal for e.g. LMMs), but there also exist rich corpora of text and multimodal data produced by players, viewers, fans, etc. Game text data, such as transcripts, have already been used to train LLM players [37, 38]. On the other hand, LLMs struggle with both spatial reasoning and planning by their very nature, while most games rely heavily on both aspects. From strategy boardgames and digital games (where long-term planning is crucial) to first-person shooters (which hinge on precision in spatial reasoning and a reactive plan for reaching the enemy base), such games remain state-of-the-art testbeds for gameplaying AI [141, 142] and will likely be fraught arenas for LLM research. Games also hinge on long-term interactions, especially in the case of LLM-based Game Masters (see Section 3.6), and thus can form testbeds or benchmarks to explore the limits of recollection under different context lengths (a critical limitation of LLMs detailed in Section 5). In terms of game design tasks, we also note that games are complex constrained problems, with hard constraints on e.g. levels that can be completed [143], but also soft constraints regarding game balance between competing players in multi-player games [144, 145], or the progression and pacing of a single-player experience [95]. While some LLMs can handle some hard constraints via, say, function calling [120], this may not be possible for more complex or more constrained game domains. Moreover, soft constraints would need to be conveyed to the LLM in more nuanced ways.
Game benchmarks specific to LLMs have already started to emerge [121], but identifying critical game-based challenges for LLMs, appropriate and interesting benchmarks, and (ethically sourced) data for training or fine-tuning LLMs remains an open question.
@@ 5 Limitations of LLMs in games
Large language models have exciting potential for games, but they also come with inherent limitations. Mainly, LLMs suffer from hallucinations [48, 146], meaning that they will output plausible but false statements simply because they are a probable sequence of words. Hallucinations are inevitable given how the world is described to the machine [147]; LLMs lack grounding, so the text they generate is detached from constraints of reality. Yet LLMs always “act” confidently in their responses, even when wholly mistaken. Beyond hallucinations, LLMs suffer from factual errors [148, 149, 150], outputting responses that are wrong even though the LLM has access to information that proves otherwise. In the context of digital games, these limitations affect certain applications of LLMs more than others, for example NPCs may hallucinate quests that do not exist in the game, or a player assistant may provide suggestions to the user based on wrong assumptions.
@@
Another limitation is that LLMs sometimes struggle to capture user intent. This is especially evident with expressions of sarcasm [151]. The ability to capture user intent is important for applications of LLMs that converse directly with the player. An LLM that correctly understands user intent then also correctly understands the context of the conversation, which is not always the case for current LLMs. Many LLMs make errors in correctly understanding user requests [152], and clarifying to the LLM multiple times results in a frustrating experience for the user. This limitation is most relevant to cases where the LLM is in direct conversation with the user, e.g. as design assistant, player assistant, or Game Master. Depending on how much the LLM output controls the user experience (e.g. as Game Master or offering production assistance to a human designer), the inability to capture user intent can be a source of frustration.
@@
On a larger scale, LLMs suffer from losing context, and struggle with continuity. This is because the “memory” of an LLM is constrained by its context size, which limits the extent of its inputs and outputs, as well as its response time due to the attention mechanism [16]. The longer the conversation, the less likely it is that the LLM will recall early events [153]. In digital games, it is possible to separately summarize the game events (see Section 3.4) and process them as part of the input to the LLM. As a game progresses past a few game sessions, however, this summary may still be too long, or details of increasing significance will be omitted, thus leading to a degraded performance. This is especially relevant for roles requiring long-term engagement, such as LLM-powered retellers or Game Masters. In Infinite Craft (see Section 3.7), this is handled by an external database that stores and looks up past combination rules, ensuring consistency in future uses of the same mechanic. However, LLMs could theoretically tackle this issue directly.
Recent models have tried to address this recollection issue by increasing the context length, with some of the larger models encompassing 128K or even 10M tokens [154]. Despite this being adequate for a wide range of applications, it may still fall short when applied to long-term tracking of game states. In particular, massive multiplayer online games offer a simulation space with a large intricate domain of actions and interactions, which scales exponentially with the number of agents (players or otherwise) participating. Researchers have also tried to address the context limit by including compressive memory into the attention mechanism of the LLM [155], in an attempt to create a seemingly infinite context length. The authors of [155], however, acknowledge its current limitation, partly due to the difficulty of selecting and compressing the data which should be “memorized”.
A Retrieval-Augmented Generation (RAG) system [156] could address this limitation, drawing from a database containing vector representations or other latent representations of pertinent text or data. When the text generator processes a sequence, the RAG system would retrieve similar entries from this external data source. This would hypothetically provide a streamlined archive of game events and actions for the LLMs to consult in order to generate a consistent narrative progression.
@@
Yet another limitation of LLMs that prevents their application in mainstream media is their cost. Running AAA games and LLMs in parallel on consumer hardware is infeasible for most players [157] due to their computation requirements. If one wants to integrate LLMs in games, they would have to host the models on their own servers or access existing models via APIs. Additionally, the cost of querying LLMs is a recurring cost, and cannot be properly estimated beforehand. This kind of problem is also affected by the scale of LLMs-powered games or tools. Similarly to how server costs increase with the number of active players in massive multiplayer online games, the more players use a LLM over multiple play sessions, the more the game developers or publishers will have to bear the financial burden. The monetary cost of this approach can be prohibitive or difficult to estimate for real-world applications. The game need not even be played by other players: to evaluate the performance of their simulations with multiple LLM-based NPCs, Park et al. ran simulations for several days with a cost of “thousands of dollars in token credits” [43]. While promising techniques to reduce the costs of running LLMs exist [158, 159], these are not yet widespread and require further engineering to set them up properly.
Perhaps due to the above limitations, the implementation and deployment of LLMs to digital game applications is still very limited. A digital game is a domain where responsiveness is vital for players, so it follows that LLMs should also be able to provide their responses quickly. Unfortunately, while research on more efficient and faster architectures is being carried out [160], the real-time application of LLMs is still not plausible. This is especially evident in other domains such as design applications, where “real time” responses are generated in around 30 seconds to over a minute [161].
@@ 6 Ethical Issues with LLMs in games
With the improvement of AI methods applied to games over the recent years, many questions regarding their ethics and real-world impact have been raised [162]. Using LLMs raises ethical issues regarding sustainability, copyright, explainability, and biases. Naturally, each of these issues has serious implications in the field of games.
@@
The reliance of LLMs on training data and training time raises concerns regarding their carbon footprint. Beyond training costs, inference over the model’s lifespan has a greater environmental impact due to constant querying [163, 164]. Factors like renewable and local energy, better model architectures, and more meaningful (and thus less wasteful) training data can mitigate this. In the context of LLMs for digital games, sustainability remains crucial, considering the carbon footprint of frequent queries during gameplay (e.g. for Game Master or NPC responses, or for LLM-powered players). This is especially pertinent if the LLM is intended to run locally, on consumer-level hardware which are usually powered by non-renewable sources.
@@
When it comes to copyright, issues apply to the input data, the output data, and the model itself. LLMs trained on data under copyright is an unfortunate common practice [165], deservedly raising public outrage [166, 167]. The models themselves have different copyright licenses applied, which can also lead to artifacts they generate to fall under the public domain [17, 168]. For the game industry, matters of IP and copyright are extremely important. This is as much a concern regarding having the company’s copyrighted content somehow used as training data by competitors, as it is about LLMs producing material that the company cannot copyright. It is important to note here that, at least when it comes to the latter concern, the role the LLM takes is very pertinent. If an LLM or LMM produces content automatically (see Section 3.8), past legal consensus in the USA indicates that the material can not be copyrighted [169]. If an LLM or LMM acts as an “assistive tool” [170] to a designer (especially for conceptual assistance, see Section 3.9) then the extensive and impactful human effort needed to transform these concepts into game design and game art likely makes the final product eligible for copyright [170]. The limited rulings in copyright courts regarding this, however, and the “likely” caveat we include in our own text, understandably would make game companies hesitant to tread in untested waters for major game IPs beyond e.g. small-scale indie productions [171]. For researchers, however, the ethical issues of copyright breach and exploitation by large corporations, and the public outcry for the above, leave a bad taste and make research in LLMs less palatable [172].
@@
In applications, understanding how a final result or product is reached is extremely crucial, particularly when a product is iteratively refined as with design assistants (see Section 3.9). This is a problem of explainability [123], whereas LLMs are inherently opaque in their generation process. Liu et al. [173] highlight different methods to improve the explainability of language models, such as concept-based explanations or saliency maps. Particularly for LLMs, the self-explanation applied via the chain-of-thought [174] reasoning has received attention by the research community [175, 176]. While this method adds a layer of explained reasoning to the generated output, there are multiple examples in the literature that demonstrate how this reasoning may just be an illusion of reasoning capabilities. Such examples include disregarding the provided reasoning in the final output [177], or reaching the correct solution via incorrect steps in math problems [178]. In the domain of games, explainability is paramount across roles, ensuring gameplay coherence and user engagement.
Replicability of an application’s behavior is equally crucial. When applying LLMs to digital games (especially as game mechanics, NPCs, or automated designers) one would expect their output quality to not change over time. This is not the case for closed-source LLMs: even when using the same model name, the same request at one time can generate content that is vastly different from a past iteration [179]. In this case, developers may have to consider switching to open-source models with open weights, such as those hosted on the popular HuggingFace Transformers library [180]. An additional benefit of switching to self-hosted models, model size notwithstanding, is the additional level of privacy that is guaranteed to the users. Querying local models ensures all messages remain within the application, whereas interacting with models hosted via APIs entails that conversations are exchanged over third-party websites. A developer might be willing to share conversation logs with an API provider [^14] Such as ChatGPT, which shares conversations by default unless opted out for model improvement. However, users may not be aware of this practice or its implications. Local deployment of LLMs has been democratized by making models more accessible on lower-end hardware, relying on the widespread adoption of the GGUF format [^15] Details of this file format are available at https://github.com/ggerganov/ggml/blob/master/docs/gguf.md and the release of different versions of the same model at varying degrees of quantization [181]. The quantization of an LLM usually results in a loss of performance, but this is usually considered a valid trade-off for the reduced model size to load on VRAM. Combined with friendly APIs for running LLMs locally, such as Open WebUI [182] and LM Studio [183], it is possible to run LLMs in a more controlled fashion. The more pertinent technological breakthroughs lie in compacting size and carbon footprint while retaining high-quality LLM outputs.
@@
Finally, biases emerge as LLMs are trained on a large corpus, usually scraped from the (Western-focused part of the) internet. This allows models to capture a current reality snapshot, which is advantageous for a conversational or question-answering model, though it requires curating this data from different kinds of biases. Some biases, such as social stereotypes, could be targeted and alleviated; others, such as exclusionary norms, pose greater challenges. In games, we identify two main concerns when interacting with an LLM: toxic behavior, and stereotypes or incorrect notions. Toxic behavior is a harmful property that a language model may learn from its training corpus, which often contains text from community-based fora or social platforms. Tools that combat toxic language in digital games are constantly evolving, with some even blocking chat messages before they are even delivered to the user [184, 185]. Therefore, similar applications could theoretically be developed to target toxic outputs from language models. Unlike human players, however, when an LLM plays the role of an NPC, it should align with the game themes and avoid any kind of toxic language or racial slurs. This requires developers to ensure proper behavior of the model through data cleaning, if the model is trained from scratch, or supplying tailored data if finetuning it to their needs. Addressing prejudices such as stereotypes and incorrect notions is complex, as they are not necessarily related to single words or expressions, but instead present themselves as a collection of ideals that can be wrong at best, and harmful at worst. An NPC LLM may exhibit real world stereotypes that can impact negatively the player experience, although we argue that the impact of prejudices from an LLM commentator or Game Master is much stronger and disturbing due to their perceived authority.